OpenStudy (anonymous):

find the domain of f(t) =6t = t^2

OpenStudy (anonymous):

Which is it?

OpenStudy (anonymous):

it is supposed to say f(t) = 6t+ t^2

Directrix (directrix):

The domain is the set of acceptable values of t. Can you think of any value of t that CANNOT be substituted for t in 6t+ t^2? Values that would not be acceptable would be numbers that make a denominator 0, a square root radicand negative; that sort of thing.

Directrix (directrix):

@CaitMcNair0404

OpenStudy (anonymous):

Thank You @Directrix, but I still do not understand because there is no denominator in my equation. I have two quizzes and 1 test due in a week, and I still have nothing done because the website has been down. I am freaking out. Would you mind helping me out with this one step by step for future references? Would 0 be a domain for this? because 6*0=0 and 0^2 would be 0 so it would mean f(t)=0. Am I going about this all wrong?

OpenStudy (anonymous):

Typically there are some implicit assumptions about these problems.

OpenStudy (anonymous):

The implicit assumption is that the function maps from real numbers to real numbers.

Directrix (directrix):

Is there any real number which could not be included in the domain?

OpenStudy (anonymous):

What do you mean @wio?

OpenStudy (anonymous):

By default, the function maps from all real numbers to all real numbers.

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

As @Directrix said, if you have 0 in denominator or <0 in a square root, then your function can't output a real number.

OpenStudy (anonymous):

So any such points would be removed from the domain.

OpenStudy (anonymous):

However in this case, nothing like that happens. This means the domains remains all real numbers.

OpenStudy (anonymous):

Yes, so a domain is the number that would make the answer invalid. Such as if it were f(x)= 2/x 0 would be a domain because it would make it 2/0. Is that correct?

OpenStudy (anonymous):

Wrong, the domain is any place where the answer is still valid.

OpenStudy (anonymous):

Though, when we say 'valid' what we mean is the result is a real number.

OpenStudy (anonymous):

so for f(x)=2/x what would be the domain? I'm trying to find an easy equation that I maybe could understand so that I could possibly "dumb down" the harder ones

OpenStudy (anonymous):

I understand that for my original question you said my domain would be all real numbers. THis is an online class and it says that my answers have to be in interval notation. my book is NO help in this. I promise I have read every word, every page, every number, every example and I still do not understand @wio .. BUT I do tremendously appreciate your help.

OpenStudy (anonymous):

The domain would be \((-\infty,0)\cup (0,\infty)\)

OpenStudy (anonymous):

Or \( x\in \mathbb R, x\neq 0\)

OpenStudy (anonymous):

The domain is expressed as a mathematical structure known as a "set"

OpenStudy (anonymous):

Even \(x\in \mathbb R \setminus \{0\}\) works.

OpenStudy (anonymous):

Interval notation generally uses \((,)\ [,]\)

OpenStudy (anonymous):

It uses \(\cup\) to combine two intervals.

OpenStudy (anonymous):

My online answer box will not let me type /.

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

(−∞,0)∪(0,∞)

OpenStudy (anonymous):

try that.

OpenStudy (anonymous):

It said incorrect. :/ it is typed exactly as you have it.

OpenStudy (anonymous):

Maybe (-infinity, 0) union (0, infinity)

OpenStudy (anonymous):

honestly, I hate how these problems are picky about input.

OpenStudy (anonymous):

when I type the word infinity, the ∞ symbol automatically appears.

OpenStudy (anonymous):

how about union?

OpenStudy (anonymous):

I hae the problems period lol.

OpenStudy (anonymous):

if union doesn't work, you could try comma ,

OpenStudy (anonymous):

Oh wait!

OpenStudy (anonymous):

The answer is \((-\infty, \infty)\) for the original problem.

OpenStudy (anonymous):

I was talking about \(3/x\) before.

OpenStudy (anonymous):

YES!!!!!!! IT WORKED!!!! THANK YOU SO MUCHHHH!!!! If it is ok with you, would I be able to request your help on here for future help?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Thank you!